OpenStudy (anonymous):

diffrentiate y=2x+1/2x-1

OpenStudy (campbell_st):

rewrite it and use the product rule \[y = (2x +1)(2x-1)^{-1}\] let u = 2x + 1 u' = 2 \[ v = (2x -1)^{-1}...... v' = -2(2x-1)^{-2}\] then \[\frac{dy}{dx} = u \times v' + v \times u'\] hope this helps